# Parameter Adjustments in POMDP-Based Trajectory Planning for Unsignalized Intersections

### v1

@@ 1 INTRODUCTION
Our contributions are twofold. At first, we verify the capability of the POMDP-based method to plan collision-free trajectories in simulations using real-life traffic data from aerial footage of two intersections with different topologies [^1] See our supplementary YouTube video.. In addition, based on those simulations, we analyze the influence of the parameter values on the method’s performance.
@@ 3 APPROACH
This section outlines the components of our approach to addressing the unsignalized intersection crossing problem. We begin by providing a formal problem definition in Section 3.1. Key to our strategy is formulating this problem as a POMDP, which we discuss in Section 3.2. To solve the POMDP, we apply an approximate solver called the Adaptive Belief Tree (ABT) algorithm, detailed in Section  3.3. Our method involves discretizing the intersection topology, a process explained in Section 3.4, where we identify all feasible 2D paths without lane changes based on combinations of intersection entrances and exits. Subsequently, we introduce dynamical and observation models in Sections 3.5, and 3.6, respectively, that capture the complexities of the real world in a simplified yet sufficiently accurate manner for trajectory planning. Entities such as state and observation are defined, and we construct a reward function in Section 3.7. This function is designed to be maximized within the POMDP, ensuring that the resulting trajectory aligns with our expectations, e.g., encompasses crucial factors such as collision avoidance and adherence to velocity limits. To complete the process, we leverage the mentioned ABT solver for our POMDP formulation to derive the desired trajectory as a sequence of longitudinal accelerations along the desired path toward the goal exit of the intersection.
@@ 3.1 Formal Problem Definition
To formally define the problem, we seek a collision-free trajectory through a known intersection topology along a predefined path $\bm{\phi}(t):t\rightarrow\mathbb{R}^{2},\,\,t\in[0,\ d]$ where $d$ is the total length of the path. We assume that the position $\bm{p}_{i}=[x_{i},\ y_{i}]^{\top}$, velocity $\bm{v}_{i}=[v_{x,i},\ v_{y,i}]^{\top}$, unit heading vector $\bm{\theta_{i}}$, width $w_{i}$, and length $l_{i}$ measurements of all $n$ relevant other (non-ego) vehicles $i\in\{1,\dots,n\}$ are available every sample time $k$. Thus, we are looking for a sequence of longitudinal accelerations $a_{0}^{k}$ of the ego vehicle (denoted with the zero subscript), such that there is no overlap between the bounding rectangles given by $\bm{p}_{0}^{k}$, $\bm{\theta}_{0}^{k}$, $w_{0}$, $l_{0}$ and $\bm{p}_{i}^{k}$, $\bm{\theta}_{i}^{k}$, $w_{i}$, $l_{i}$ for each $i$. Lateral accelerations of all vehicles need not to be directly determined, as the curvature of the followed path implicitly controls it, provided that drifting is avoided.
@@ 3.2 POMDPs
Fundamentally, $T$ represents the stochastic dynamics model of the system, and $Z$ represents the stochastic observation (measurement) model with the probabilities $T(s^{\prime}|a,s)$ of transition from state $s\in S$ to state $s^{\prime}\in S$ with action $a\in A$, and $Z(o|a,s^{\prime})$ of observation $o\in O$ in $s^{\prime}$. The particular way of modeling the dynamical and measurement properties for our purposes is described Sections 3.5 and 3.6.
Since the current state is unknown, the belief ${b\in B}$ is introduced, representing a distribution over all states from the set $B$ known as the belief space:
@@
where $\propto$ denotes that the right-hand side probabilities are not normalized. This converts the original POMDP problem into an MDP where beliefs are states, with continuous and observable state space B.
The goal of the POMDP planner is to determine a policy $\pi:B\rightarrow A$, that maximize the expected sum of discounted rewards called the value function $V$:
@@ 3.3 Adaptive Belief Tree Algorithm
The ABT algorithm [Kurniawati and Yadav, 2016] is an online and anytime POMDP solver that uses the Augmented Belief Tree for finding well-performing policies. The tree $\mathcal{T}$ is constructed by sampling initial belief $b_{0}$ with $n_{\mathrm{par}}$ particles and propagating them with actions according to the observation and dynamics models $T$ and $Z$ until the set depth of the optimization horizon $N$ or a terminal state is reached. At each step, the particles are assigned a reward $r=R(s,a)$.
Sequences of quadruples $(s_{i},a_{i},o_{i},r_{i})$ representing a path in the tree where $i$ is the depth of each contained node are sampled state trajectories also referred to as episodes. At last, each episode is assigned a quadruple $(s_{N},-,-,r_{N})$ where $s_{N}$ is a sampled next state and $r_{N}$ is the expected future reward for the following states. This is computed directly for the terminal states; non-terminal states are heuristically estimated, in our case, with a 3-step lookahead. An example of an Augmented Belief Tree is illustrated in Figure 2.
@@
With the tree containing $n_{\mathrm{ep}}$ episodes, the policy based on a $Q(b,a)$ (Q-value) estimate $\hat{Q}(b,a)$ for each belief-action pair is obtained:
@@
where $Q(b,a)$ is the reward for performing an action $a$ in the first step and then behaving optimally. The estimate is computed as an average reward of all episodes $h\in H(b,a)$ containing the pair $(b,a)$ in depth $l$:
@@
The current belief is updated after the action $\pi(b_{0})$ is applied, and new measurements are obtained. The corresponding branches of the belief tree are reused, with the current belief being the new root of the tree. This process is repeated until the goal is reached. If needed, new particles are generated to maintain the $n_{\mathrm{par}}$ number of particles to avoid particle depletion.
@@ 3.4 Topology Discretization
To identify possible paths, we utilize the lanelet2 [Poggenhans et al., 2018] map in the OSM XML format as a source for the full intersection topology. Using the lanelet2 library, we generate a route graph, searching for all possible paths as a sequence of $\mathbb{R}^{2}$ points. Subsequently, each path is stored as two separate cubic splines for the $x$ and $y$ coordinates. This provides an efficient mapping from the path position to the global coordinates, denoted as $\bm{\phi}(p):p\rightarrow\mathbb{R}^{2}$, and mapping to the unit heading $\bm{\psi}(p):p\rightarrow\bm{y}\in\mathbb{R}^{2}:\|\bm{y}\|=1$. For the inverse mapping converting global coordinates back to the position on a path, denoted as $\epsilon(\bm{p}):\bm{p}\rightarrow\mathbb{R}$, we employ a search algorithm to determine the nearest point on the selected path.
@@ 3.5 Dynamics Model
As we proceed with the formulation of our POMDP, it is crucial to introduce a dynamical model describing how the state, defined in the previous section, evolves over time. While various mathematical models of different complexities exist for vehicles, we aim for simplicity and reduced computational demand. Thus, we follow the approach from [Hubmann et al., 2018] and adopt a discrete linear point mass model for each vehicle $i\in\{0,\dots,n\}$:
@@
$\bm{A}=\begin{bmatrix}1&\Delta t&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\qquad\bm{B}=\begin{bmatrix}\frac{1}{2}{\Delta t}^{2}\\ \Delta t\\ 0\end{bmatrix}\,.$ (8)
Here, $\Delta t$ represents the time between subsequent samples $k$ and $k+1$, and $u(\bm{s})$ denotes the acceleration of the vehicle. The Gaussian noise $\bm{\eta}$ with zero mean and covariance matrix $\bm{Q}$ accounts for model inaccuracies.
@@
We implemented the IDM model for its simplicity:
@@
The acceleration of the ego vehicle is obtained from the resulting POMDP policy according to (4). The model (7), and (8) together with other vehicles acceleration predictions (9) represent the dynamics model referred to as $T$ in Section 3.2.
@@ 3.6 Observation Model
Figure 4: Likelihood functions of features
@@
where $\mu_{i}^{k}$ signifies the specific path for the mappings, and $\bm{\zeta}$ represents Gaussian observation noise with zero mean and covariance matrix $\bm{R}$.
@@ 3.7 Reward Function
The reward function is a crucial component that defines the behavior of the ego vehicle within the intersection. It is responsible for aligning the vehicle’s actions with our desired objectives: efficient navigation through the intersection while prioritizing safety, avoiding excessive acceleration, and preventing collisions. This multi-goal task is formulated as
@@
where $R_{\mathrm{vel}}$ is a positive weight. The desired velocity $v_{\textrm{des}}$ is influenced by factors such as the curvature $\kappa_{\mu}(p_{0}^{k})$ at the vehicle’s position $p_{0}^{k}$ along its intended path $\mu_{0}^{k}$. The desired velocity is determined as:
@@
where $v_{\textrm{curve}}=\sqrt{a_{\textrm{lat},\textrm{max}}/\kappa_{\mu}(p_{0}^{k})}$ is the velocity limited by the maximal acceptable lateral acceleration $a_{\textrm{lat},\textrm{max}},\textrm{max}$, and $v_{\textrm{lim}}$ maximal velocity limit of the intersection.
@@ 4 RESULTS
In this section, we test the investigated method’s collision-free planning capabilities and the impact of parameter adjustments on the planning algorithm’s effectiveness. For these proposes, we ran simulations based on data collected from aerial observations of two specific intersections as shown in Figure 5. In the simulations, building upon the results of [Maňour, 2021], we replaced a yielding vehicle with our ego vehicle to mimic real-life situations. Even though the use of such offline data does not include direct reactions of other vehicles to the ego vehicle, we deem it appropriate for testing situations where the ego is expected to yield and adequately respond to the behavior of others. Also, our simulations assume that the ego vehicle adheres to the point mass model (7).
@@ 4.1 Roundabout Scenario
The results, depicted in Figures 6 and 7, show the mean of normalized reward $\Delta t\sum_{i}{r^{i}}$ achieved for various settings. Settings without crashes had values between $-3200$ and $-3600$. Mean reward below $-4000$ represents occurring crashes. The reward is also based on $\Delta t$, as the smaller the time step, the higher the reward because the planning algorithm has more samples to react.
@@ 4.2 Threeway Junction Scenario
Note that the reward here does not indicate the total quality of the solution; rather, it says how well the vehicle follows the reference velocity. Slowing down information gathering is reflected as a lower reward.
@@ 5 CONCLUSION
ACKNOWLEDGEMENTS
This work was co-funded by the European Union under the project ROBOPROX (reg. no. CZ.02.01.0100/22_0080004590) and by the Technology Agency of the Czech Republic under the project Certicar CK03000033.

### v2

@@ 1 INTRODUCTION
Our contributions are twofold. At first, we verify the capability of the POMDP-based method to plan collision-free trajectories in simulations using real-life traffic data from aerial footage of two intersections with different topologies [^1] See: video link unavailable for double-blind review. In addition, based on those simulations, we analyze the influence of the parameter values on the method’s performance.
@@ 3 APPROACH
This section outlines the key components of our approach to addressing the unsignalized intersection crossing problem. We begin by providing a formal problem definition in Section 3.1. Key to our strategy is formulating this problem as a POMDP, which we discuss in Section 3.2. To solve the POMDP, we apply an approximate solver called the Adaptive Belief Tree (ABT) algorithm, detailed in Section  3.3. Our method involves discretizing the intersection topology, a process explained in Section 3.4, where we identify all feasible 2D paths without lane changes based on combinations of intersection entrances and exits. Subsequently, we introduce dynamical and observation models in Sections 3.5, and  3.6, respectively, that capture the complexities of the real world in a simplified yet sufficiently accurate manner for trajectory planning. Essential entities such as state and observation are defined, and we construct a reward function in Section 3.7. This function is designed to be maximized within the POMDP, ensuring that the resulting trajectory aligns with our expectations, e.g., encompasses crucial factors such as collision avoidance and adherence to velocity limits. To complete the process, we leverage the already mentioned ABT solver for our POMDP formulation to derive the desired trajectory as a sequence of longitudinal accelerations along the desired path toward the goal exit of the intersection.
@@ 3.1 Formal Problem Definition
To formally define the problem, we seek a collision-free trajectory through a known intersection topology along a predefined path $\bm{\phi}(t):t\rightarrow\mathbb{R}^{2},\,\,t\in[0,\ d]$ where $d$ is the total length of the path. We assume that the position $\bm{p}_{i}=[x_{i},\ y_{i}]^{\top}$, velocity $\bm{v}_{i}=[v_{x,i},\ v_{y,i}]^{\top}$, unit heading vector $\bm{\theta_{i}}$, width $w_{i}$, and length $l_{i}$ measurements of all $n$ relevant other (non-ego) vehicles $i\in\{1,\dots,n\}$ are available every sample time $k$. Thus, we are looking for a sequence of longitudinal accelerations $a_{0}^{k}$ of the ego vehicle (denoted with the zero subscript), such that there is no overlap between the bounding rectangles given by $\bm{p}_{0}^{k}$, $\bm{\theta}_{0}^{k}$, $w_{0}$, $l_{0}$ and $\bm{p}_{i}^{k}$, $\bm{\theta}_{i}^{k}$, $w_{i}$, $l_{i}$ for each $i$. Lateral accelerations of all vehicles need not be directly determined, as the curvature of the followed path implicitly controls it, provided that drifting is avoided.
@@ 3.2 POMDPs
Fundamentally, $T$ represents the stochastic dynamics model of the system, and $Z$ represents the stochastic observation (measurement) model with the probabilities $T(s^{\prime}|a,s)$ of transition from state $s\in S$ to state $s^{\prime}\in S$ with action $a\in A$, and $Z(o|a,s^{\prime})$ of observation $o\in O$ in $s^{\prime}$. The particular way of modeling the dynamical and measurement properties for our purposes is described Sections 3.5 (dynamical model) and 3.6 (observation model).
Since the current state is unknown, the so-called belief $b\in B$ is introduced, representing a distribution over all states from the set $B$ known as the belief space:
@@
where the $\propto$ sign denotes that the right-hand side probabilities are not normalized. This converts the original POMDP problem into an MDP where the beliefs are the states, with the continuous fully observable state space B.
The goal of the POMDP planner is to determine a policy $\pi:B\rightarrow A$, that should maximize the expected sum of discounted rewards called the value function $V$:
@@ 3.3 Adaptive Belief Tree Algorithm
The ABT algorithm [Kurniawati and Yadav, 2016] is an online and anytime POMDP solver that uses the Augmented Belief Tree for finding well-performing policies. The tree $\mathcal{T}$ is constructed by sampling initial belief $b_{0}$ with $n_{\mathrm{par}}$ particles and propagating them with actions according to the observation and dynamics models $T$ and $Z$ until the set depth of the optimization horizon $N$ or a terminal state is reached. At each step, the particles are also assigned a reward $r=R(s,a)$.
Sequences of quadruples $(s_{i},a_{i},o_{i},r_{i})$ representing a path in the tree with $i$ being the depth of each contained node are sampled state trajectories also referred to episodes. At last, each episode is assigned an additional quadruple $(s_{N},-,-,r_{N})$ where $s_{N}$ is a sampled next state and $r_{N}$ is the expected future reward for all following states. This is computed directly for the terminal states; non-terminal states are heuristically estimated, in our case, with a 3-step lookahead. An example of an Augmented Belief Tree is illustrated in Figure 2.
@@
With the tree containing $n_{\mathrm{ep}}$ episodes, the policy based on a Q-value estimate $\hat{Q}(b,a)$ for each belief-action pair is obtained:
@@
where Q-value $Q(b,a)$ is a reward for performing an action $a$ in the first step and then behaving optimally. The $\hat{Q}(b,a)$ estimate is computed as an average reward of all episodes $h\in H(b,a)$ containing the pair $(b,a)$ in depth $l$:
@@
The current belief is updated after the action $\pi(b_{0})$ is applied, and new measurements are obtained. The corresponding branches of the belief tree are reused, with the current belief being the new root of the tree. The whole process is repeated until the goal is reached. If needed, new particles are generated to maintain the $n_{\mathrm{par}}$ number of particles to avoid particle depletion.
@@ 3.4 Topology Discretization
To identify possible paths, we utilize the lanelet2 [Poggenhans et al., 2018] map in the OSM XML format as a source for the full intersection topology. Using the lanelet2 library, we generate a route graph, searching for all possible paths as a sequence of $\mathbb{R}^{2}$ points. Subsequently, each path is stored as two separate cubic splines for the $x$ and $y$ coordinates. This setup provides an efficient mapping from the path position to the global map coordinates, denoted as $\bm{\phi}(p):p\rightarrow\mathbb{R}^{2}$, and mapping to the unit heading $\bm{\psi}(p):p\rightarrow\bm{y}\in\mathbb{R}^{2}:\|\bm{y}\|=1$. For the inverse mapping converting global coordinates back to the position on a path, denoted as $\epsilon(\bm{p}):\bm{p}\rightarrow\mathbb{R}$, we employ a search algorithm to determine the nearest point on the selected path.
@@ 3.5 Dynamics Model
As we proceed with the formulation of our POMDP, it is crucial to introduce a dynamical model describing how the state, defined in the previous section, evolves over time. While various mathematical models of different complexities exist for vehicles, we aim for simplicity and reduced computational demand. Thus, we follow the approach from [Hubmann et al., 2018] and adopt a discrete linear point mass model for each vehicle $i\in\{1,\dots,n\}$:
@@
$\bm{A}$ $=\begin{bmatrix}1&\Delta t&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\qquad\bm{B}=\begin{bmatrix}\frac{1}{2}{\Delta t}^{2}\\ \Delta t\\ 0\end{bmatrix}\,.$
Here, $\Delta t$ represents the time between subsequent samples $k$ and $k+1$, and $u(\bm{s})$ denotes the acceleration of the vehicle. The additive Gaussian noise $\bm{\eta}$ with zero mean and covariance matrix $\bm{Q}$ accounts for model inaccuracies.
@@
We implemented the IDM-based model for its simplicity:
@@
The acceleration of the ego vehicle is the input of our POMDP, obtained from the resulting policy according to (4). The model (7) together with other vehicles acceleration predictions (8) represent the POMDP dynamics model referred to as $T$ in Section 3.2.
@@ 3.6 Observation Model
Refer to captionRefer to caption
Figure 4: Likelihood functions of features– $f_{1}$ (left), $f_{2}$ (right).
@@
where $\mu_{i}^{k}$ signifies the specific path for the mappings, and $\bm{\zeta}$ represents observation noise following a Gaussian distribution with zero mean and covariance matrix $\bm{R}$.
@@ 3.7 Reward Function
The reward function is a crucial component that defines the behavior of the ego vehicle within the intersection. It is responsible for aligning the vehicle’s actions with our desired objectives: efficient navigation through the intersection while prioritizing safety, avoiding excessive acceleration, and preventing collisions. This multi-goal task is mathematically formulated as
@@
where $R_{\mathrm{vel}}$ is a positive weight. The desired velocity $v_{\textrm{des}}$ is influenced by factors such as the curvature $\kappa_{\mu}(p_{0}^{k})$ at the vehicle’s position $p_{0}^{k}$ along its intended path $\mu_{0}^{k}$. The desired velocity is determined as follows:
@@
where $v_{\textrm{curve}}=\sqrt{a_{\textrm{lat},\textrm{max}}/\kappa_{\mu}(p_{0}^{k})}$ is the velocity limited by the maximal acceptable lateral acceleration during cornering, and $v_{\textrm{lim}}$ is the velocity speed limit of the intersection.
@@ 4 RESULTS
In this section, we test the investigated method’s collision-free planning capabilities and the impact of parameter adjustments on the planning algorithm’s effectiveness. For these proposes, we ran simulations based on data collected from aerial observations of two specific intersections as shown in Figure 5. In the simulations, building upon the results of [Author-Removed, 2021] we replaced a yielding vehicle with our ego vehicle to mimic real-life situations. Even though the use of such offline data does not include direct reactions of other vehicles to the ego vehicle, we deem it appropriate for testing situations where the ego is expected to yield and adequately respond to the behavior of others. Also, our simulations assume that the ego vehicle adheres to the point mass model (7).
@@ 4.1 Roundabout Scenario
The results, depicted in Figures 6 and 9, show the mean of normalized reward $\Delta t\sum_{i}{r^{i}}$ achieved for various settings. In most cases, the simulation resulted in no crashes with reward values between $-3200$ and $3600$ based on $\Delta t$. The smaller the time step, the higher the reward because the planning algorithm has more samples to react.
@@ 4.2 Threeway Junction Scenario
Note that the reward here does not indicate the total quality of the solution; rather, it says how well the vehicle follows the reference velocity. Slowing down to gather more information is reflected as a lower reward.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption